# On Root Capacity, Intersection Indicium, Minimal Generating Sets of Galois Closure & Compositum Feasible Triplets

### v2

@@ Abstract.
We carry forward the work started by the author and Bhagwat in [2] and develop the Theory of root clusters further in this article and also apply similar methods to resolve certain problems in related areas. We establish the Inverse root capacity problem for number fields which is a generalization of Inverse cluster size problem for number fields proved in [2]. We introduce the concept of intersection indicium as a generalization of the concept of ascending index introduced in [2] and prove some of its properties. We then establish the Inverse intersection indicium problem for number fields (excluding certain cases) which is a generalization of Inverse ascending index problem for number fields proved in [2]. We give a field theoretic formulation for the concept of minimal generating sets of splitting fields which was introduced by the author and Vanchinathan in [10] and generalize a result in [10] for number fields and also establish the existence of field extensions over number fields for given degree and given cardinality of minimal generating set of Galois closure dividing the degree. We generalize a result in [6] by establishing that a certain family of triplets is compositum feasible over any number field and we also list all the irreducible triplets in this family. We also prove a partial case of a conjecture in [6]. In the concluding section of this article, we improve on the inverse problems proved in [2] and this article by proving that there exist arbitrarily large finite families of pairwise non-isomorphic extensions having additional properties that satisfy the given conditions.
@@ 1. Introduction
The Theory of root clusters was substantially developed by the author and Bhagwat in their work in [2] which built on previous work by Perlis in [15] and Krithika and Vanchinathan in [11]. This article is yet another contribution in enriching the theory further by means of generalizations and we also apply similar methods to resolve certain problems in related areas.
Let $K$ be a perfect field. We fix an algebraic closure $\bar{K}$ once and for all and work with finite extensions of $K$ contained in $\bar{K}$. Let $L/K$ be a degree $n$ extension and $\tilde{L}$ be its Galois closure inside $\bar{K}$. Let $G={{\rm Gal}}(\tilde{L}/K)$ and $H={{\rm Gal}}(\tilde{L}/L)$. We have the notion of cluster size of $L/K$, $r_{K}(L)$ which is $N_{G}(H):H]$ (See Section 2.1 in [2] for basic properties of cluster size of field extension). From Section 3.2 in [2], number of clusters of $L/K$, $s_{K}(L)$ is $G:N_{G}(H)]$ which is also the number of distinct fields inside $\bar{K}$ isomorphic to $L$ over $K$.
In Section 2, we prove some interesting properties of unique intermediate extensions for given extensions. This notion was introduced in Section 7 of [2]. The concepts of strong cluster magnification and root capacity were introduced by the author and Bhagwat in Sections 4 and 6 of [2] respectively. In Section 3, we establish the Root Capacity Magnification Theorem, Theorem 3.6 which is a generalization of Cluster Magnification Theorem proved in [11]. The concept of cluster towers was introduced in [11]. We give a field theoretic formulation for cluster towers and prove Proposition 3.8 about strong cluster magnification and cluster towers.
@@
The notion of minimal generating sets of the splitting field of a polynomial was introduced by the author and Vanchinathan in Section 2 in [10]. In Section 7, we give a field theoretic formulation for minimal generating sets and prove Proposition 7.4 about strong cluster magnification and minimal generating sets. We prove Theorem 7.6 as a generalization of Theorem 2.11 in [10] for number fields. We then go on to establish Theorem 7.7 (by applying similar methods as in earlier sections) which is as follows. For notations see Section 7.
@@
Then we prove Proposition 8.10 which is a partial case of Conjecture 4 in [6] where it is conjectured that the set of compositum feasible triplets over $K$ is a commutative multiplicative monoid. The notion of a compositum feasible triplet being irreducible was introduced in [13]. In Proposition 8.14, we list all irreducible triplets in the family of compositum feasible triplets covered in Theorem 8.7. Propositions 8.12 and 8.15 cover some other interesting cases of compositum feasible triplets and sum feasible triplets.
In Section 9, we improve on the inverse problems proved in [2] and this article by proving that there exist arbitrarily large finite families of pairwise non-isomorphic extensions having additional properties that satisfy the given conditions. The following is Theorem 9.2 which is an improvement on Inverse cluster size problem for number fields Theorem 3.1.1 in [2].
@@
We obtain similar improvements for Theorem 9.0.5 in [2] and Theorems 5.2, 7.6, 7.7 and 8.7 which are Theorems 9.3, 9.4, 9.5, 9.6 and 9.7 respectively.
@@ 2. Some Remarks on Unique Intermediate Extensions
Proposition 2.2 helps us to give alternate proofs for certain cases in Section 7.3 in [2] which is encapsulated as the following example.
@@ 5. Inverse Root Capacity Problem
Let $K$ be a number field. Let $n>2$ and $r|n$. Then there exists an extension $L/K$ of degree $n$ with cluster size $r$ with Galois closure having solvable Galois group.
@@ 6. Inverse Intersection Indicium Problem
Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$ with Galois closure having solvable Galois group.
@@
Let $K$ be a number field and $n>2$ and $t|n$ with $t>1$. Let $r=n/t$ for $t<n$ and $r=n$ for $t=n$. By the proof of Theorem 6.1 in [2], we have $L/K$ with $[L:K]=n$ and $t_{K}(L)=t$ and $r_{K}(L)=r$ which is the same as the field extension $L/K$ in second approach of proof of Theorem 5.2. Consider the notations as in that proof above.
@@
    If $K$ is a number field then for any $n\in\mathbb{N}$, we have a degree $n$ extension $L^{\prime}/K$ such that $Gal(\tilde{L^{\prime}}/K)=\mathfrak{S}_{n}$ and $L^{\prime}\cap L=K$.
@@
    The proof is similar to proof of Part (a) of Main theorem in [12] but we provide it nonetheless. By Lemma 5.3, there exist infinitely many degree $n$ extensions over $K$ with Galois closures having Galois group ${\mathfrak{S}}_{n}$.
    Since $L/K$ is separable, it has finitely many intermediate fields. Thus there is a degree $n$ extension $L^{\prime}/K$ such that $L^{\prime}\not\subset L$ and $Gal(\tilde{L^{\prime}}/K)=\mathfrak{S}_{n}$. Now $L^{\prime}\cap L$ is an intermediate field for $L^{\prime}/K$. Since $\mathfrak{S}_{n-1}$ is a maximal subgroup of $\mathfrak{S}_{n}$, we have $L^{\prime}\cap L=K$ or $L^{\prime}\cap L=L^{\prime}$. The latter case is not possible since $L^{\prime}\not\subset L$. Hence $L^{\prime}\cap L=K$.
@@ 7. Minimal Generating Sets of Galois Closure
Suppose $M/K$ is obtained by strong cluster magnification from $L/K$ through $F/K$. Then $B=\{L_{i_{j}}\}_{j=1}^{m}\subset S$ is a minimal generating set of the Galois closure of $L/K$ if and only if $B^{\prime}=\{L_{i_{j}}F\}_{j=1}^{m}$ is a minimal generating set of the Galois closure of $M/K$.
@@
Let $K$ be a number field and $n>2$ an odd composite number. Then there exists an $L/K$ of degree $n$ for which the Galois closure has minimal generating sets of cardinalities $2,3,\dots,\omega(n)$ (where $\omega(n)$ is the number of distinct prime divisors of $n$) and these are the only possible cardinalities for minimal generating set of $\tilde{L}$.
@@
By Theorem 2.11 in [10] we have the result for $K=\mathbb{Q}$ that is we have $L^{\prime}/\mathbb{Q}$ satisfying the required properties. Now Galois group of Galois closure of $L^{\prime}/\mathbb{Q}$ is $G=\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ which is solvable. Also $H=Gal(\tilde{L^{\prime}}/L^{\prime})=1\times(\mathbb{Z}/n\mathbb{Z})^{\times}$.
By Lemma 6.6 (2), $G$ is realizable as a Galois group over $K$ say for $E/K$. Let $L/K$ be fixed field of $H$ inside $E/K$. By Remark 7.1.1, $L/K$ satisfies the required properties. ∎
@@
    Let $B$ be any minimal generating set of the Galois closure of $L/K$. Then $B\supset\{L_{i}\ |\ L_{S\backslash L_{i}}\neq\tilde{L}\}$.
@@
    Suppose $B\subset S$ is the unique minimal generating set of the Galois closure of $L/K$. Then $B=\{L_{i}\ |\ L_{S\backslash L_{i}}\neq\tilde{L}\}$.
@@
    Let $L_{i}$ be such that $L_{S\backslash L_{i}}\neq\tilde{L}$. Assume on the contrary that $L_{i}\not\in B$. Thus $B\subset S\backslash L_{i}$. Thus $L_{B}\subset L_{S\backslash L_{i}}$. Hence $L_{B}\neq\tilde{L}$ which is a contradiction.
@@
    Consider $L_{i}$ such that $S\backslash L_{i}=\tilde{L}$. Now since $S\backslash L_{i}$ is a generating set of $\tilde{L}$, by Lemma 7.2 we have a minimal generating set $B^{\prime}$ of $\tilde{L}$ such that $B^{\prime}\subset S\backslash L_{i}$. By uniqueness of $B$ we have $B=B^{\prime}$. Thus $B\subset S\backslash L_{i}$. Hence $L_{i}\not\in B$. By part (1) we are done.

### v3

# arXiv:2505.20238v3[math.NT] 28 Feb 2026

@@ Abstract.
We carry on the work started by the author and Bhagwat and develop the theory of root clusters further in this article and also resolve certain problems in related areas. We introduce some new notions as well as recall earlier notions for field extensions over a perfect base field: root cluster size, its generalization root capacity, its dual notion ascending index and its generalization intersection indicium. We establish our results on the Inverse problems for these generalized notions over number fields which generalizes our earlier results. We give a field theoretic formulation for the concept of minimal generating sets of splitting fields which was introduced by the author and Vanchinathan. We present new results as well as generalizations of our earlier results on the cardinalities of minimal generating sets for extensions over number fields. We generalize a result of Drungilas et al. by establishing that a certain family of triplets is compositum feasible over any number field and we also list all the irreducible triplets in this family. We also prove a partial case of a conjecture of Drungilas et al. We also improve on the inverse problems by proving that there exist arbitrarily large finite families of pairwise non-isomorphic extensions having additional properties that satisfy the given conditions. Our methods for all these problems are Galois theoretic in nature and heavily rely on the known cases of the inverse Galois problem.
@@ 1. Introduction
The theory of root clusters was substantially developed by the author and Bhagwat in their work in [1] which builds on previous work by Perlis in [15] and Krithika and Vanchinathan in [11]. This article is yet another contribution in enriching the theory further by means of systematic generalizations to more expansive frameworks that provide metrics that allow for a more nuanced measurement of how roots of irreducible polynomials are distributed within various field extensions. We also apply similar methods to resolve certain problems in related areas.
Let $K$ be a perfect field. We fix an algebraic closure $\bar{K}$ once and for all, and work with finite extensions of $K$ contained in $\bar{K}$. Let $L/K$ be a degree $n$ extension and let $\tilde{L}$ be its Galois closure inside $\bar{K}$. Let $G={{\rm Gal}}(\tilde{L}/K)$ and $H={{\rm Gal}}(\tilde{L}/L)$. We have the notion of cluster size of $L/K$, $r_{K}(L)$ which is $N_{G}(H):H]$ (See Section 2.1 in [1] for basic properties of cluster size of field extension). From Section 3.2 in [1], number of clusters of $L/K$, $s_{K}(L)$ is $G:N_{G}(H)]$ which is also the number of distinct fields inside $\bar{K}$ isomorphic to $L$ over $K$.
In Section 2, we prove some interesting properties of unique intermediate extensions for given extensions. This notion was introduced in Section 7 of [1]. The concepts of strong cluster magnification and root capacity were introduced by the author and Bhagwat in Sections 4 and 6 of [1] respectively. In Section 3, we recall these notions and establish the Root Capacity Magnification Theorem, Theorem 3.6 which is a generalization of Cluster Magnification Theorem proved in [11]. The concept of cluster towers was introduced in [11]. We give a field theoretic formulation for cluster towers and prove Proposition 3.8 about strong cluster magnification and cluster towers.
@@
The notion of minimal generating sets of the splitting field of a polynomial was introduced by the author and Vanchinathan in Section 2 in [10]. In Section 7, we give a field theoretic formulation for minimal generating sets and prove Proposition 7.4 about strong cluster magnification and minimal generating sets. We prove Theorem 7.6 as a generalization of Theorem 1.1 in [10] for number fields. We also prove Theorem 7.8 as a generalization of Theorem 1.2 in [10]. We then go on to establish Theorem 7.9 (by applying similar methods as in earlier sections) which is as follows. For notations see Section 7.
@@
Then we prove Proposition 8.11 which is a partial case of Conjecture 4 in [6] where it is conjectured that the set of compositum feasible triplets over $K$ is a commutative multiplicative monoid. The notion of a compositum feasible triplet being irreducible was introduced in [13]. In Proposition 8.15, we list all irreducible triplets in the family of compositum feasible triplets covered in Theorem 8.7. Propositions 8.13 and 8.16 cover some other interesting cases of compositum feasible triplets and sum feasible triplets.
We also improve on the inverse problems proved in [1] and this article by proving that there exist arbitrarily large finite families of pairwise non-isomorphic extensions having additional properties that satisfy the given conditions. The following is Theorem 5.4 which is an improvement on Inverse cluster size problem for number fields Theorem 3.1.1 in [1].
@@
We obtain similar improvements for Theorem 9.0.5 in [1] and Theorems 5.5, 7.6, 7.9 and 8.7 which are Theorems 6.2, 5.8, 7.7, 7.10 and 8.8 respectively.
@@ 2. Some Remarks on Unique Intermediate Extensions
Proposition 2.2 helps us to give alternate proofs for certain cases in Section 7.3 in [1]. The following example encapsulates those cases with the alternate proofs.
@@ 5. Inverse Root Capacity Problem
Let $K$ be a number field. Let $n>2$ and $r|n$. Then there exists an extension $L/K$ of degree $n$ with cluster size $r$.
Lemma 5.2 (Lemma 3.1.4, [1]).
Lemma 5.3.
Proof.
The following is an improvement on Theorem 5.1.
Theorem 5.4.
Proof.
Remark 5.4.1.
Remark 5.4.2.
@@
An improvement on Inverse root capacity problem for number fields Theorem 5.5 (Similar improvement as in Theorem 5.4).
Theorem 5.8.
@@ 6. Inverse Intersection Indicium Problem
Let $K$ be a number field. Let $n>2$ and $t|n$. Then there exists an extension $L/K$ of degree $n$ with ascending index $t_{K}(L)=t$.
An improvement on Theorem 6.1 (Similar improvement as in Theorem 5.4).
Theorem 6.2.
@@
Let $K$ be a number field and $n>2$ and $t|n$ with $t>1$. Let $r=n/t$ for $t<n$ and $r=n$ for $t=n$. By the proof of Theorem 6.1 in [1], we have existence of an extension $L/K$ with $[L:K]=n$ and $t_{K}(L)=t$ and $r_{K}(L)=r$ which is the same as the field extension $L/K$ in second approach of proof of Theorem 5.5. Consider the notations as in that proof above.
@@
    If $K$ is a number field then for any $n\in\mathbb{N}$, we have an extension $L^{\prime}/K$ such that $Gal(L^{\prime}/K)=\mathfrak{S}_{n}$ and $L^{\prime}\cap L=K$.
@@
    The proof is similar to proof of Part (a) of Main theorem in [12] but we provide it nonetheless. By Lemma 5.6, there are infinitely many pairwise linearly disjoint extension $L^{\prime}/K$ such that $Gal(L^{\prime}/K)=\mathfrak{S}_{n}$. Call the infinite set of these extensions as $\mathfrak{S}$.
    Since $L/K$ is separable, it has finitely many intermediate extensions. Consider the subset $\mathfrak{S}^{\prime}$ of $\mathfrak{S}$ consisiting of those $L^{\prime}/K$ which contain atleast one non-trivial intermediate extension of $L/K$. Due to pairwise linear disjointness, $\mathfrak{S}^{\prime}$ is finite. Thus we are done.
@@ 7. Minimal Generating Sets of Galois Closure
Remark 7.1.2.
Suppose $B=\{L_{i_{j}}\}_{j=1}^{m}\subset S$ is a minimal generating set of the Galois closure of $L/K$. It is easy to verify that for any $\sigma\in Gal(\tilde{L}/K)$ we have that $B^{\prime}=\sigma(B)=\{\sigma(L_{i_{j}})\}_{j=1}^{m}\subset S$ is also a minimal generating set of $\tilde{L}$.
@@
Suppose $M/K$ is obtained by strong cluster magnification from $L/K$ through $F/K$. Then $B=\{L_{i_{j}}\}_{j=1}^{m}\subset S$ is a minimal generating set of the Galois closure of $L/K$ if and only if $B^{\prime}=\{L_{i_{j}}F\}_{j=1}^{m}$ is a minimal generating set of the Galois closure of $M/K$. In this case $|B|=|B^{\prime}|$.
@@
Let $K$ be a number field and $n$ be an integer such that $\frac{n}{2^{v_{2}(n)}}>2$ is composite (where $v_{2}$ is usual $2$-adic valuation). Then there exists an $L/K$ of degree $n$ for which the Galois closure has minimal generating sets of cardinalities $2,3,\dots,\omega^{\prime}(n)$ (where $\omega^{\prime}(n)$ is the number of distinct odd prime divisors of $n$) and these are the only possible cardinalities for minimal generating set of $\tilde{L}/K$.
@@
Let $n^{\prime}=\frac{n}{2^{v_{2}(n)}}$. Thus $n^{\prime}>2$ is an odd composite number. By Theorem 1.1 in [10], there exists an $L^{\prime}/\mathbb{Q}$ of degree $n^{\prime}$ for which the Galois closure has minimal generating sets of cardinalities $2,3,\dots,\omega(n^{\prime})$ (where $\omega(n^{\prime})$ is the number of distinct prime divisors of $n^{\prime}$) and these are the only possible cardinalities for minimal generating set of $\tilde{L^{\prime}}$. Observe that $\omega^{\prime}(n)=\omega(n^{\prime})$.
Now from the proof in [10], we have that the Galois group of Galois closure of $L^{\prime}/\mathbb{Q}$ is $G=\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$ which is solvable. Also $H=Gal(\tilde{L^{\prime}}/L^{\prime})=1\times(\mathbb{Z}/n\mathbb{Z})^{\times}$. By Lemma 6.6 (2), $G$ is realizable as a Galois group over $K$ say for $E/K$. Let $M/K$ be fixed field of $H$ inside $E/K$. By Remark 7.1.1, $M/K$ has degree $n^{\prime}$ and its Galois closure has minimal generating sets of cardinalities $2,3,\dots,\omega(n^{\prime})$ and these are the only possible cardinalities for minimal generating set of $\tilde{M}/K$.
By Lemma 5.7, there exists a cyclic Galois extension $F/K$ of degree $2^{v_{2}(n)}$ such that $\tilde{M}$ and $F$ are linearly disjoint over $K$. Let $L=MF$. By Theorem 3.3, $L:K]=n^{\prime}\cdot 2^{v_{2}(n)}=n$. By Proposition 7.4, the Galois closure of $L/K$ has minimal generating sets of cardinalities $2,3,\dots,\omega^{\prime}(n)$ and these are the only possible cardinalities for minimal generating set of $\tilde{L}/K$. Also observe that $E=\tilde{M}$ and $\tilde{L}=\tilde{M}F=EF$ and $Gal(\tilde{L}/K)\cong Gal(E/K)\times Gal(F/K)\cong G\times(\mathbb{Z}/(2^{v_{2}(n)})\mathbb{Z})$ which is solvable. ∎
An improvement on Theorem 7.6 (Similar improvement as in Theorem 5.4).
Theorem 7.7.
Let $K$ be a number field and $n$ be an integer such that $\frac{n}{2^{v_{2}(n)}}>2$ is composite. Then we get arbitrarily large finite families of degree $n$ extensions $L/K$ inside $\bar{K}$ which are pairwise non-isomorphic over $K$ and are pairwise linearly disjoint over $K$, for which the Galois closure has minimal generating sets of cardinalities $2,3,\dots,\omega^{\prime}(n)$ and these are the only possible cardinalities for minimal generating set of Galois closure of $L/K$.
The following is a generalization of Theorem 1.2 in [10].
Theorem 7.8.
Let $K$ be a number field and $n>2$ be an integer.
  1. (1)
    For any $n>2$ and $d\mid n$ with $d>2$, there exists an $L/K$ of degree $n$ for which the Galois closure has all its minimal generating sets of cardinality $k$ (and all cluster towers of length k+1) for the following values of $k$ : (i) k=d1k=d-1, (ii) k=d2k=d-2, (iii) $k=2$.
  2. (2)
    For $K=\mathbb{Q}$, there exists an $L/\mathbb{Q}$ of degree $n$ for which the Galois closure has all its minimal generating sets of cardinality $k$ (and all cluster towers of length k+1) for the following values of $n$ and $k$:
    1. (a)
      (i) $n$ is a multiple of 12 and $k=5$, (ii) $n$ is a multiple of 11 and $k=4$.
    2. (b)
      $n$ is a multiple of $p+1$ where $p$ is an odd prime and $k=3$.
Proof.
  1. (1)
    Consider any number field $K$ and any $n>2$ and $d\mid n$ with $d>2$. By Theorem 1.2 (1) in [10], there exists an $L^{\prime}/K$ of degree $d$ for which the Galois closure has all its minimal generating sets of cardinality $k$ for the following values of $k$ : (i) k=d1k=d-1, (ii) k=d2k=d-2. By Lemma 5.7, there exists a Galois extension $F/K$ of degree $n/d$ such that $\tilde{L^{\prime}}$ and $F$ are linearly disjoint over $K$. Let $L=L^{\prime}F$. By Theorem 3.3, $L:K]=d\cdot(n/d)=n$. By Proposition 7.4, the Galois closure of $L/K$ has all its minimal generating sets of cardinality $k$.
    Now consider any number field $K$ and any $n>2$ and $k=2$. Let $q>2$ be a prime power such that $q\mid n$. By Theorem 1.2 (2) in [10], there exists an $L^{\prime}/\mathbb{Q}$ of degree $q$ for which the Galois closure has all its minimal generating sets of cardinality $2$. Now from the proof in [10], we have that the Galois group of Galois closure of $L^{\prime}/\mathbb{Q}$ is $G=\mathbb{F}_{q}\rtimes(\mathbb{F}_{q})^{\times}$ which is solvable. Let $H=Gal(\tilde{L^{\prime}}/L^{\prime})$. By Lemma 6.6 (2), $G$ is realizable as a Galois group over $K$ say for $E/K$. Let $M/K$ be fixed field of $H$ inside $E/K$. By Remark 7.1.1, $M/K$ has degree $q$ and its Galois closure has all its minimal generating sets of cardinality $2$. By Lemma 5.7, there exists a Galois extension $F/K$ of degree $n/q$ such that $\tilde{M}$ and $F$ are linearly disjoint over $K$. Let $L=MF$. Thus $[L:K]=n$ and by Proposition 7.4, the Galois closure of $L/K$ has all its minimal generating sets of cardinality $2$.
  2. (2)
    By Theorem 1.2 (2) in [10], there exists an $L^{\prime}/\mathbb{Q}$ of degree $n^{\prime}$ for which the Galois closure has all its minimal generating sets of cardinality $k$ for the following values of $n^{\prime}$ and $k$:
    1. (a)
      (i) $n^{\prime}=12$ and $k=5$, (ii) $n^{\prime}=11$ and $k=4$.
    2. (b)
      $n^{\prime}=p+1$ where $p$ is an odd prime and $k=3$.
    Now the result follows by applying Lemma 5.7 and Proposition 7.4.
We establish an interesting result now.
@@
An improvement on Theorem 7.9 (Similar improvement as in Theorem 5.4).
Theorem 7.10.
@@
    Let $B$ be any minimal generating set of the Galois closure of $L/K$. Then $B\supset D$.
@@
    We have that either $D=\emptyset$ or $D=S$. In the second case $S$ is the only minimal generating set of the Galois closure of $L/K$.
  3. (3)
    Suppose $B\subset S$ is the unique minimal generating set of the Galois closure of $L/K$. Then $B=S$.
@@
[…]
@@
[…]
  3. (3)
[…]